A machine of efficiency 80% is used to raise a load of a body of mass 76kg through a vertical height of 3m in 30s. Calculate the power input (g=10ms²)
1
Expert's answer
2020-12-14T12:13:51-0500

In order to raise a load the machine should deliver the following power:


"P_d = \dfrac{mgh}{t}"

where "m = 76kg" is the mass of the load, "g = 10m\/s^2" is the gravitational acceleration, "h = 3m" is the vertical height, and "t = 30s" is the time of raising. Since the machine is 80% efficient, The input power should be:


"P_i = \dfrac{P_d}{0.8} = 1.25 \dfrac{mgh}{t}\\
P_i = \dfrac{76\cdot 10\cdot 3}{30} = 76\space W"

Answer. 76 W.